### v2

# 1Huawei Technologies, 2NTU, 3HKU, 4CUHK

Abstract
[…]
(a) Performance comparisons.
(b) Performance breakdown.
Figure 1: Overview of the performance by SWE-Lego models and its breakdown analysis. (a) SWE-Lego models establish a new frontier on SWE-bench Verified, outperforming same-scale competitors. Notably, our results are based on hack-free evaluation, whereas prior works scores could be inflated by the Git hacking. (b) Our hybrid SWE-Lego dataset delivers the largest boost of +25.6%; subsequent refined SFT adds +3.8%, and TTS contributes +6.2%, together lifting the Qwen3-32B model from 23.2% to 58.8%.
1 Introduction
Recent advances of large language models (LLMs) have led to remarkable progress in autonomous software engineering (SWE) agents, capable of repository-scale issue resolving (Yang et al., 2024; Zhang et al., 2024; Xia et al., 2024). Unlike short, single-file tasks (Du et al., 2023; Austin et al., 2021; Chen et al., 2021a; Li et al., 2022), SWE agents are able to navigate code repositories, faithfully reproduce failing tests, identify buggy files and lines, implement fixes, and validate within executable sandboxes (Yang et al., 2023; Liu et al., 2024; Shrivastava et al., 2023), which demands long-horizon reasoning, robust tool use, and multi-turn interaction abilities.
To equip LLMs with these capabilities, a common approach is supervised fine-tuning (SFT), which distills knowledge from stronger SWE agents. However, prior efforts often lack datasets that include execution environments and realistic bug instances needed to produce high-quality, diverse trajectories (Wang et al., 2025b; Yang et al., 2025c; Jain et al., 2025). Beyond SFT, performance on SWE tasks can be further improved via mid-training, but this typically requires substantially more computation and data (Copet et al., 2025; Yang et al., 2025d). Alternatively, recent work applies reinforcement learning (RL) without a fixed teacher model or predefined SWE trajectories (Luo et al., 2025; Sonwane et al., 2025; Wei et al., 2025), yet limited executable instances remain a bottleneck. In addition, RL is more computationally intensive, demands heavier training infrastructure, and is prone to training collapse due to hyperparameter sensitivity. These practical constraints motivate a critical question: how far can we push a lightweight, SFT-only training recipe for SWE tasks?
To investigate, we propose SWE-Lego, a simple yet powerful SFT framework for training SWE agents. We discover that with careful data design, training techniques, and test-time scaling for inference, SFT alone can achieve state-of-the-art performance, matching or surpassing more complex training paradigms under similar LLM scales, as shown in Figure 1. SWE-Lego comprises three core building blocks:
  • We introduce the SWE-Lego dataset, a collection of 32k high-quality task instances and 18k validated expert trajectories. The SWE-Lego dataset is constructed from two complementary sources: high-quality real-world GitHub pull requests (Badertdinov et al., 2025), and scalable synthetic instances (Yang et al., 2025c), where both sources complement each other in quality and quantity: the real-world data provides the majority of benefits but is hard to scale, while the synthetic data provides additional gains with further scaling. Furthermore, we implement rigorous trajectory validation and filtering, such as preventing agents from Git hacking (i.e., inspecting Git history to recover solutions) to ensure the data teaches genuine problem-solving skills.
  • We refine the conventional supervised fine-tuning to better train SWE agents. While standard SFT approaches have been widely applied for training (Yang et al., 2025c; Wang et al., 2025b), they typically treat every token in an expert trajectory equally during loss calculation. This overlooks a critical issue: trajectories often contain both successful actions and intermediate errors, and naively learning from mistakes could reinforce undesired behavior. To address this, we introduce step-level error masking, which excludes tokens associated with execution-time errors (e.g., tool failures or failing tests) from the loss. Moreover, we introduce another refinement of difficulty-based curriculum learning, which trains the model on easier tasks before progressing to harder ones, with difficulty estimated empirically by trajectory length.
  • To trade additional compute for further gains, we study test-time scaling (TTS)  (Jain et al., 2025; OpenHands Team, 2025; Luo et al., 2025) along two orthogonal dimensions: (i) sequential scaling by increasing the maximum number of interaction turns, and (ii) parallel scaling via multiple rollouts with verifier selection. We seek to explore the best TTS strategies for SWE tasks given the compute constraints. We find that sequential scaling saturates around 100–140 turns; beyond this point, compute is better allocated to parallel rollouts. In addition, generative verifiers (Pan et al., 2025; Jain et al., 2025) consistently outperform regression-based scorers (OpenHands Team, 2025) for parallel scaling.
Empirically, SWE-Lego establishes new state-of-the-art results among open-source models on SWE-bench Verified (Jimenez et al., 2024): SWE-Lego-Qwen3-8B reaches 42.2%, and SWE-Lego-Qwen3-32B attains 52.6% using only SFT (Figure 1(a)). With test-time scaling, the performance rises to 49.6% and 58.8% under TTS@16, respectively. These results demonstrate that a meticulously crafted SFT pipeline can rival or exceed the performance of more complex and computationally expensive training regimes. Notably, our results are obtained without Git hacking, whereas prior reports are often inflated by such solution leakage. Ablations in Figure 1(b) show that our SWE-Lego dataset yields a substantial boost of +25.6%, refined SFT contributes an additional +3.8%, and TTS ultimately adds +6.2%, which together lift Qwen3-32B from 23.2% to 58.8%, confirming the effectiveness of each building blocks. We hope SWE-Lego can provide a reproducible and lightweight post-training paradigm for building effective SWE agents, and the associated data and models of this work will be open-sourced for future research in the community.
2 SWE-Lego Dataset: Combining Real-world and Synthetic Data
2.1 Overview
LLMs are increasingly integrated into autonomous agents for real-world SWE tasks. We adopt the SWE-Agent paradigm (Yang et al., 2024), in which an LLM operates through a specialized agent-computer interface that allows it to autonomously navigate repositories, search and edit code, and execute tests to diagnose and resolve issues. In this context, a SWE task instance is usually defined as the concrete SWE issue within a codebase, comprising a natural-language problem description, test cases, a reference golden patch, and related artifacts. Besides, a trajectory is the complete sequence of actions taken by an agent and the interleaved observations from environments. To train capable SWE agents, our primary goal is to curate a large-scale, diverse, and executable collection of task instances, together with high-quality, validated trajectories.
Dataset Type Executable Repositories Task Instances Valid Trajectories
R2E-Gym-Train Synthetic Yes 10 4.6k 3.3k
SWE-Swiss Synthetic Yes <<600 10k NA
SWE-Fixer Real No 856 115k NA
SWE-Gym Real Yes 11 2.4k 491
SWE-bench-Train Real No 37 19k NA
SWE-rebench Real Yes 3.5k 21k NA
SWE-smith Real & Synthetic Yes 128 50k 5k
SWE-Lego Real & Synthetic Yes 3251 32.1k 14.1k resolved + 4.0k semi-resolved
Table 1: Comparison of public SWE issue-resolving datasets and our SWE-Lego dataset. The proposed dataset combines real and synthetic instances at scale, with executable environments and a large pool of validated trajectories, providing a stronger foundation for training SWE agents. Note that SWE-Swiss and SWE-Fixer release only sub-trajectories, and therefore, counts of full trajectories are not reported. “Semi-resolved” means the trajectories do not pass all test cases but have partial correctness, e.g., correct file localization.
Figure 2: Our SWE‑Lego pipeline comprises three stages: environment construction from over 3,000 repositories; hybrid task creation by combining real pull requests with synthetic bugs; and expert‑trajectory generation and curation for SFT.
(a) Scaling of Valid Trajectories
(b) Performance Scaling on SWE-bench Verified
Figure 3: Impact of synthetic data augmentation on data scale and performance. (a) The number of valid expert trajectories increases substantially with the addition of synthetic instances, demonstrating an effective way to expand supervision beyond limited real-world PRs. (b) The resolve rate on SWE-bench Verified rises as more synthetic instances are added per repository, across different repository counts, indicating that hybrid data improves not only dataset scale but also downstream model effectiveness. “Insts./Repo” denotes the average number of instances per repository.
We adopts a hybrid data construction strategy that combines real-world and synthetic SWE task instances, as shown in Figure 2. This is motivated by the empirical observation that neither data source alone is sufficient: real-world instances, while authentic, are inherently limited in quantity given strict filtering criteria. On the other hand, synthetic instances, while scalable, lack the complexity of natural software repositories (Table 3). We therefore mix the real with synthetic instances, and apply rigorous generation and validation procedures to produce high-quality trainable trajectories.
The resulting SWE-Lego dataset comprises 32k executable task instances and 18k validated expert trajectories to date, as summarized in Table 1. Compared with existing open-source SWE datasets, SWE-Lego spans more repositories, offers more executable task instances, and contains substantially more validated trajectories. Moreover, scaling this hybrid corpus consistently yields better downstream performance on SWE-bench Verified, as evidenced in Figure 3. In the following subsections, we detail the pipeline used to construct the SWE-Lego dataset.
2.2 SWE Task Instance Creation and Validation
Repository Collection and Sandbox Construction.
The proposed dataset is built upon SWE-rebench (Badertdinov et al., 2025), a large collection of real-world executable repositories. We select over 3,000 Python-centric repositories with permissive licenses and active maintenance histories. To ensure reproducibility, we deploy a fully automated pipeline that parses configuration files (e.g., setup.py) to build Docker containers. Only repositories that successfully build and pass sanity tests are retained, serving as the verifiable “base plates” for our subsequent task construction.
Real-world and Synthetic Task Construction.
Real-world tasks are derived from resolved GitHub pull requests (PRs) paired with corresponding issue descriptions. They offer high authenticity with production-level bug complexity, but are often labor-intensive and limited in quantity due to the strict filtering criteria (Badertdinov et al., 2025; Pan et al., 2025; Jain et al., 2025; Zeng et al., 2025). Moreover, each real-world task requires a unique sandbox aligned with the pre-fix snapshot. While SWE-smith (Yang et al., 2025c) proposes a “PR mirror” method to generate real-world instances (prompting LLMs to revert PR edits via .diff plaintext instead of checking out the base commits), we find this approach yields lower-quality data compared to directly collecting PRs that satisfy SWE-rebench (Badertdinov et al., 2025)’s filtering rules. Thus, we adopt the latter strategy for real-world task construction.
Synthetic tasks, by contrast, are generated via active bug injection following SWE-smith (Yang et al., 2025c), leveraging two core techniques: (i) LLM Rewrite: prompting models to rewrite code using only function headers and docstrings; (ii) AST Reformulation: extracting abstract syntax trees (ASTs) for classes/functions and applying random transformations, e.g., removing conditionals/loops, modifying operators or dependencies. Synthetic tasks enable high scalability and efficiency, as they are not constrained by the availability of filtered real PRs. Even repositories with no qualified PRs can be used to create large-scale tasks at low cost. Additionally, synthetic tasks from the same repository share a single sandbox environment, since all bugs are injected into the same codebase version.
These two data sources are complementary for scaling SWE tasks: real-world data provides depth (complexity and realism), while synthetic data provides breadth (quantity and coverage). Table 2 summarizes the core-field differences between the two, and additional implementation details are provided in Appendix A.1. Table 3 presents a quantitative comparison of complexity. We observe a stark contrast: real-world instances touch more files and change more lines than synthetic ones. This confirms that real-world data captures the sprawl of complex bugs, whereas synthetic data offers focused, surgical training signals and yields a higher rate of valid trajectories for training. Further analysis of issue categories is provided in Appendix A.2. Figure 3 demonstrates the effectiveness of this hybrid approach: for a fixed set of repositories, scaling synthetic data, e.g., from “real only” to “real + 5 synthetic insts./repo”, consistently improves both the number of valid expert trajectories and the resolve rate of the trained models. Together with Table 1, this shows that scaling executable hybrid data, rather than real instance alone, is key to obtaining richer supervision signals and higher downstream performance.
Core Field Real-World Tasks Synthetic Tasks
golden_patch Merged PR diff as the ground-truth fix. Inverse the patch of the programmatically injected bug.
FAIL-TO-PASS & PASS-TO-PASS Label of test sets that induced by running the original project tests before and after the merged patch. Label of test sets that induced by running the same tests on buggy vs. restored (pre-injection) code.
problem_statement Original GitHub issue linked to the PR. LLM-generated description conditioned on the golden patch, test files, and logs.
image_name (Sandbox) Per-PR sandbox tied to the exact pre-merge commit. Shared sandbox for many injected bugs on the same base commit given the repository.
Table 2: High-level comparison of how core fields are instantiated for real-world and synthetic SWE tasks in SWE-Lego. Both sources populate the same unified schema, but differ in data origin and sandbox granularity, leading to realistic yet costly real instances and scalable, controllable synthetic ones.
Category Metric Real-World Tasks Synthetic Tasks Real-World + Synthetic Tasks
Instances 18409 13710 32119
Issue Length Words 135 168 149
Golden Patch Files 3.7 1.0 2.6
Hunks 9.5 1.3 6.0
Lines 138.0 18.8 87.3
Test Cases Fail-to-Pass 12.4 10.9 11.7
Pass-to-Pass 74.6 41.2 60.4
Total 87.0 52.1 72.1
Valid Trajectories 5010 9100 14110
Table 3: Characteristic comparisons of task instances. Real-world tasks are more complex, touching more files and modifying more lines, while synthetic tasks are more focused and yield more valid, fully resolving trajectories. The combined dataset covers a broader spectrum of difficulty.
2.3 Trajectory Rollout and Validation
With the hybrid task instances, we now roll out expert trajectories that are later used for SFT. We adopt the OpenHands scaffold (v0.53) (Wang et al., 2024) with Qwen3-Coder-480B-A35B-Instruct (Yang et al., 2025a) as the teacher agent, and set the maximum number of interaction turns to 100. The model offers competitive coding and reasoning capabilities on challenging SWE tasks and, being open-weight, can be deployed locally at scale for trajectory distillation. During distillation, we found that the agent sometimes peeked at Git history or future commits, thereby cheating rather than faithfully resolving the issue. As showcased in Figure 4, we also observed avoidable tool-calling errors and tools that were ineffective in practice. To enhance trajectory quality, we propose and implement the following three practices.
Figure 4: Examples of problematic commands or tool interactions: (a) high‑risk commands that can cause Git hacking; (b) view_range parameter mis-specification; (c) ineffective task_tracker.
Preventing Git Hacking. Recently it is known by the SWE community that LLM agents can unexpectedly “hack” git metadata to locate the golden patch by directly inspecting commit logs [^1] https://github.com/SWE-bench/SWE-bench/issues/465, posted on Sep 3, 2025. . A demonstrative example is shown Figure 4(a), and its associated full trajectory is listed in Appendix D. To prevent such Git hacking, we sanitize the repository history exposed to the agent. For real instances, we remove all commits and log messages dated after the issue creation date so that future fixes are invisible. For synthetic instances, since the buggy version of repository is ahead from the non-buggy version due to intentional bug injection, we remove the entire git history and all logs, exposing only a single snapshot of the buggy codebase. This forces the agent to genuinely reason about code and tests rather than reading answers from version control. As expected, we see in Table 5 that the valid-trajectory rate drops slightly from an inflated 31% to 30%, and the average number of turns increases a bit from 61.8 to 62.5. In addition, as mentioned in Section 5, we also report the results of SWE-bench Verified without Git hacking throughout this paper, unless specified otherwise.
Handling Malformed Tool Errors. With Qwen3-Coder-480B-A35B-Instruct, we observe frequent malformed calls to the str_replace_editor tool, e.g., passing a string to view_range or specifying out-of-bounds line ranges (see Figure 4(b)), causing tool failures and a waste of interaction budget. To mitigate these errors, we apply lightweight post-hoc correction, i.e., parsing strings to integers and clipping ranges to valid spans, which enables the agent to inspect code more reliably. Without robust tool handling, the valid-trajectory rate decreases slightly from 30% to 29%, and the average number of turns increases from 62.5 to 66.6 (see Table 5).
Pruning Ineffective Tools. Recent work (OpenAI, 2025a; Anthropic, 2025b) has begun integrating task-management tools such as task_tracker for long-horizon planning and progress tracking. However, we find that Qwen3-Coder-480B-A35B-Instruct struggles to use them effectively, which often cause execution errors (see Figure 4(c)). We therefore discard this tool and restrict the tool set to four essential operations: execute_bash, str_replace_editor, think, and finish, to keep trajectories streamlined, as evidenced by fewer interaction turns (from 65.5 to 62.5) and a higher valid-trajectory rate (from 27% to 30%) in Table 5.
Validation and Filtering.
Quality control is enforced through rigorous post-hoc validation. We classify trajectories as resolved if they pass all tests without regression. However, not all resolved trajectories are equal, nor are all unresolved ones useless. As shown in Table 5, our filtering strategy is two-fold: (i) we filter low-quality resolved trajectories (e.g., those that “cheat” by modifying tests), which yields a modest performance gain; (ii) we recycle semi-resolved trajectories, i.e., those that correctly locate the buggy file (with the recall rate of 100%) but fail to fix it. It results 4k additional trajectories for training. Adding these trajectories provides a clear boost (+1.2%), demonstrating that even partial successes contain valuable supervision for fault localization.
Practice Avg Turns Valid Traj. Rate (%)
All 62.5 30
w/o preventing Git hacking 61.8 31
w/o handling malformed tool errors 66.6 29
w/o pruning ineffective tools 65.5 27
Table 4: Ablation of rollout practices. The valid trajectory rate is evaluated across 100 SWE-Lego instances.
Data Strategy # Traj Resolve Rate (%)
All resolved 14.6k 40.4
- low-quality resolved 14.1k 41.0
+ semi-resolved 18.1k 42.2
Table 5: Ablation of filtering strategies on SWE-bench Verified.
Takeaways We introduce the SWE-Lego dataset, comprising 32k high-quality software engineering task instances and 18k expert trajectories, the core building brick for training effective SWE agents.
We demonstrate that combining real-world and synthetic data yields complementary benefits in both quantity and quality. Moverover, trajectory curation and filtering further enhance data utility, thereby improving the issue resolving rate.
3 Refined Supervised Fine-tuning
Conventional supervised fine-tuning on expert trajectories has proven effective for SWE tasks, however, the intrinsic characteristics of these trajectories remains underexplored. A trajectory may ultimately resolve an issue, yet it often contains erroneous intermediate steps, e.g., failed attempts and malformed tool calls. Rather than blindly learning from the entire trajectory, we propose step-level error masking, which enables the model to learn from correct actions while excluding incorrect ones. Furthermore, we investigate curriculum learning in the SWE context, wherein progressively increasing trajectory difficulty better guides the training process.
3.1 Step-level Error Masking
While the high-quality trajectories identified in Section 2.3 offer valuable learning signals at a macro level, we propose a more granular refinement: step-level error masking. This approach is motivated by the observation that expert demonstrations rarely proceed monotonically to success; they often contain intermediate missteps, such as invalid function calls or incorrect arguments, that are later identified and corrected. Our method maintains the full trajectory context but selectively masks the loss calculation on erroneous agent responses. As illustrated in Figure 5, we use regular expression to identify error messages provided by the terminal environment, and apply error masking to the corresponding agent response. Crucially, errors stemming from reproducing bugs or executing test files are excluded from this masking process. This technique applies gradient updates solely to valid actions, thereby providing a robust learning signal that enhances both action generation and error recovery capabilities. By emphasizing learning from correct actions, it directly reduces core reasoning failures, such as “Incorrect Implementation” and “Localization Error”, that are prevalent in later training stages (Section 3.4). Empirical results in Table 7 demonstrate that step-level error masking improves model performance by over 2 points.
Figure 5: An example of step-level error masking, which maintains the complete trajectory context while selectively masking the loss calculation on incorrect agent responses.
3.2 Difficulty-Based Curriculum Learning
To maximize sample efficiency and training performance, we implement a curriculum learning strategy that progressively exposes the model to tasks of increasing complexity. This approach encourages the model to acquire fundamental skills from simpler examples before advancing to more challenging problems, thereby promoting more robust generalization. We explore two principled methods for difficulty categorization:
  • Model-Based Scoring: We fine-tune a Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) model as a difficulty scorer on a dataset of 1.5k human-annotated (task, difficulty) pairs. This scorer achieves 70% accuracy on a held-out test set and classifies tasks into three tiers: Easy (estimated solving time << 15 minutes), Medium (15 minutes to 1 hour), and Hard (>> 1 hour).
  • Trajectory-Length-Based Heuristic: We discover a strong negative correlation (r=0.95r=-0.95) between instance resolve rate and the number of turns in the trajectory, as depicted in Figure 7. Leveraging this finding, we partition the data into three difficulty bins based on turn count: Easy (0-50 turns), Medium (50-70 turns), and Hard (70-100 turns).
For curriculum learning, we select the trajectory-length-based heuristic because (i) it outperformed the model-based approach by 0.5 points, and (ii) it is more computationally efficient and less prone to mislabeling.
We adopt a three-stage SFT curriculum based on difficulty rankings. In each stage, the model is trained exclusively on data from the corresponding difficulty tier. To mitigate catastrophic forgetting of previously acquired skills, the training data for each subsequent stage is augmented with all data from the preceding stages. This curriculum aligns with the training dynamics observed in Section 3.4: it first grounds the model on “Easy” tasks to overcome basic “Failed to Reproduce” errors, then introduces “Hard” tasks to develop the strategic planning needed to avoid “Ran Out of Max Turns” failures. Table 7 shows that curriculum learning consistently enhances the performance of both the 8B and 32B models, with the highest results achieved when combined with error masking.
Figure 6: Correlation between number of turns and average resolve rate.
Error Masking Curriculum Learning Resolve Rate (%)
SWE-Lego-8B SWE-Lego-32B
42.2 52.6
41.8 (-0.4) 51.8 (-0.8)
40.2 (-2.0) 50.4 (-2.2)
39.4 (-2.8) 48.8 (-3.8)
Figure 7: Ablation of training strategies.
3.3 Training Details
For training, we use the Qwen3-8B/32B [^2] https://huggingface.co/Qwen/Qwen3-8B; https://huggingface.co/Qwen/Qwen3-32B (Yang et al., 2025a) series as the base models. We perform full-parameter SFT on the collected SWE-Lego trajectories using LLaMA-Factory framework (Zheng et al., 2024). Models are trained for 4 epochs with a global batch size of 64. We use the AdamW optimizer (Loshchilov and Hutter, 2019) with a weight decay of 0.01 and a cosine learning rate schedule with a warmup ratio of 0.1. The learning rate is set to 1e-4 for the 7B/8B models and 5e-5 for 32B models. To enable training with long-horizon trajectories, we use a maximum context length of 128k tokens. For the Qwen3 series, RoPE scaling techniques such as YaRN (Peng et al., 2023) are adopted to support context lengths of up to 128k tokens.
3.4 Error Analysis during Training
To validate our methodology, we analyze the agent’s failure modes during training by categorizing all failed instances (including unresolved, error, and empty patch cases) into five error types based on the agent’s trajectory:
  • Failed to Reproduce: The agent either did not attempt reproduction or attempted but failed to reproduce the issue.
  • Read Localization Error: The agent reproduces the issue but fails to open or inspect all files modified in the ground-truth (golden) patch.
  • Write Localization Error: The agent opens all required files but fails to identify the correct lines to edit, either no edits are made to ground-truth files or the edited lines have no overlap with the ground-truth patch.
  • Ran Out of Max Turns: The agent successfully reproduced the issue and correctly localized the read and write locations, but reached the maximum iteration limit (100 turns) before completing the task.
  • Incorrect Implementation: The case does not fall into any of the above categories; the agent appears to find the right location but implements an incorrect fix, resulting in failing tests.
Figure 8 shows a clear progression. Early on, “Failed to Reproduce” errors (38.97%) dominate, indicating basic task misalignment. After epoch 1, this error vanishes, but “Ran Out of Max Turns” errors spike (35.14%), revealing a strategic planning deficit. In later epochs, “Incorrect Implementation” and “Localization Error” become the bottlenecks. This shift from process failures to reasoning failures motivates our combined approach: curriculum learning sequentially addresses the early and mid-stage errors by structuring task complexity, while step-level error masking specifically hones the fine-grained reasoning needed to overcome the persistent late-stage errors.
(a) Resolve Rate Improvement Trend
(b) Error Type Distribution Across Epochs
Figure 8: Evolution of model performance and failure modes over 4 epochs. (a) Resolve rate increases steadily, reaching 42.2% by epoch 4. (b) Error modes shift from early “Failed to Reproduce” to mid “Ran Out of Max Turns” to late “Incorrect Implementation” and “Localization Error” bottlenecks.
Takeaways We refine conventional SFT for SWE tasks with two innovations: (1) step-level error masking, which enables the model to learn from effective intermediate actions, and (2) curriculum learning, which progressively increases task difficulty, approximated by the number of interaction turns.
Our refined SFT outperforms conventional SFT by 3.8%, establishing new state-of-the-art performance among open-source models of comparable size on SWE-Bench Verified: SWE-Lego-Qwen3-8B hits 42.2%, and SWE-Lego-Qwen3-32B reaches 52.6%.
4 Test-time Scaling
Test-time scaling (TTS) can effectively improve the SWE agent performance by allocating additional compute during inference. Typically there are two complementary dimensions: sequential scaling with more interaction turns, and parallel scaling that generates multiple rollouts and selects the best via a verifier. However, it still remains unclear their optimal balance, given limited constraints on compute budget or latency. In Section 4.1, we first systematically investigate this trade-off and identify when to prioritize each dimension. We further investigate the design of verifier for better parallel scaling in Section 4.2.
4.1 Balancing Sequential and Parallel Scaling
We investigate resource allocation between sequential and parallel scaling by varying both the maximum interaction turns and the number of candidate rollouts. We use latency as the cost metric for budget control. For sequential turns, as the trajectory lengthens, the growing context requires more computation per turn, resulting in latency that scales super-linearly (between linear and quadratic) with respect to the number of interaction turns; In contrast, candidate rollouts are executed in batch, so latency exhibits sub-linear scaling with respect to the number of rollouts.
As shown in Figure 9, sequential scaling is highly efficient in the low-turn regime: additional turns yields more environment feedback, enables error correction, and allows iterative refinement. This makes sequential scaling the preferred strategy under tight latency budgets. However, performance starts to saturate around 100–140 turns. The agent either succeeds within the allotted turns or encounters obstacles that additional turns alone cannot overcome (more analysis in Appendix B.1). Beyond this saturation point, parallel scaling coupled with verifier-based selection becomes more effective, as independent trajectories explore diverse paths through the solution space. The iso-latency contours in Figure 9 highlight this balance: under equivalent latency, the optimal allocation shifts from sequential-dominated to parallel-dominated as the total latency budget increases.
Figure 9: Resolve rate on SWE-bench Verified as a function of maximum interaction turns (x-axis) and number of parallel rollouts (y-axis). Curves indicate iso-latency contours where different sequential-parallel combinations yield equivalent wall-clock time. Sequential scaling is most effective at low turn counts; after saturation, improvements are driven primarily by parallel scaling.
4.2 Improved Parallel Scaling
During parallel scaling, a verifier takes a trajectory and its proposed patch and returns a score between 0 and 1 that reflects the likelihood of successful resolution. We compare two paradigms: the regressive verifier (OpenHands Team, 2025), which appends a scoring head to the backbone and is trained with binary cross-entropy loss, and the generative verifier (Pan et al., 2025; Jain et al., 2025), which formulates verification as text generation by predicting “yes” or “no” and computing the score from normalized token probabilities. The generative formulation aligns with the pre-trained next-token prediction objective, potentially better leveraging the model’s inherent knowledge (Zhang et al., 2025). For each test instance, we generate $K$ candidate rollouts and let the verifier select the top-scoring one. TTS@$K$ is the fraction of instances resolved under this top-1 selection from $K$ candidates.
(a) Generative vs. Regressive Verifier
(b) Comparison with Existing Verifiers
Figure 10: Parallel TTS performance on SWE-bench Verified. (a) Ablation on the verifier paradigm: generative verifiers consistently outperform regressive counterparts across model sizes (8B and 32B). (b) Comparison with publicly available verifiers: SWE-Lego-Verifier-8B outperforms OpenHands-Critic-32B and R2E-Gym-Verifier-14B, with the performance gap widening as rollouts increase.
Regressive vs. Generative Verifier.
We compare the two paradigms using the same training data (5k resolved, 13k unresolved trajectories from real-world tasks) and the same backbone (e.g., Qwen3-8B). Emperically, we find that larger verifiers benefit stronger rollout models, so for 3232B rollout models we use Qwen3-Coder-30B-A3B as the verifier (see Appendix B.2 for more details). For consistency with prior work, we set the maximum interaction turns to 100 in all verifier experiments. As shown in Figure 10(a), the generative verifier consistently outperforms the regressive one. At small $K$, both achieve similar TTS@$K$; as $K$ increases, the regressive verifier plateaus while the generative verifier continues to improve. On SWE-Lego-Qwen3-8B, the gap reaches 2.8% at K=16K{=}16 (49.6% vs. 46.8%). We hypothesize that it is because the generative formulation aligns with next-token prediction, better leveraging pre-trained semantics. We therefore adopt the generative paradigm.
Comparison with Existing Verifiers.
We compare SWE-Lego-Verifier-8B against publicly available verifiers on SWE-Lego-Qwen3-8B rollouts (Figure 10(b)). Our verifier achieves TTS@16 of 49.6%, outperforming OpenHands-Critic-32B (44.0%) (OpenHands Team, 2025) and R2E-Gym-Verifier-14B (47.0%) (Jain et al., 2025). Beyond absolute performance, we observe qualitatively different scaling behaviors across verifier paradigms. OpenHands-Critic-32B, which employs the regressive paradigm, exhibits performance degradation at higher $K$, a counterintuitive result indicating that larger candidate pools overwhelm its discriminative capacity. In contrast, both generative verifiers (ours and R2E-Gym) maintain monotonic improvement toward the Pass@$K$ upper bound, further confirming that the generative formulation provides more robust scaling properties. The performance advantage over R2E-Gym-Verifier-14B likely stems from two factors: a larger volume of high-quality training trajectories and better distribution alignment between training data and SWE-Lego agent behaviors.
Takeaways Sequential-then-Parallel: We suggest to prioritize sequential scaling before a certain saturation point, and then allocate remaining compute to parallel scaling.
Generative >> Regressive: The generative verifier consistently outperforms the regressive variant for parallel scaling across different model sizes and rollout budgets.
5 Comparisons with Existing Methods
We compare the performance of SWE-Lego models with both proprietary and open-source baselines on SWE-bench Verified. As mentioned in Section 2.2, Docker images of SWE-bench Verified released prior to Sep 3, 2025 may contain the ground truth commits, which can lead to Git hacking, e.g., via git log --all or by grepping commit messages. For fair comparisons with prior methods, we report our results both with and without git hacking. The results without hacking are obtained using the latest SWE-bench Verified Docker images. From Table 6, we see that our SWE-Lego-Qwen3-8B achieves a $42.2\%$ resolve rate with refined SFT and reaches 49.6% with further TTS@16, while SWE-Lego-Qwen3-32B attains $52.6\%$ with SFT and 58.8% with TTS@16. These hack-free results surpass most open-source models and several larger proprietary models, even when their results are generally inflated by git hacking.
Model Scaffold Training Resolve Rate (%)
Proprietary Models
   OpenAI-GPT-4o (OpenAI, 2024) Internal pipeline - 33.2
   OpenAI-o3 (OpenAI, 2025c) Mini-SWE-agent - 58.4
   Claude-4-Sonnet (Anthropic, 2025a) SWE-agent - 66.6
   Claude-4.5-Sonnet (Anthropic, 2025b) Mini-SWE-agent - 70.6
   OpenAI-GPT-5 (OpenAI, 2025b) OpenHands - 71.8
   Gemini 3 Pro Preview (Google, 2025) Mini-SWE-agent - 74.2
Open-Source Models
Parameters $\approx$ 7B
   Qwen2.5-Coder-Instruct-7B (Hui et al., 2024) MOpenHands - 1.0
   Qwen3-8B (Yang et al., 2025a) OpenHands - 7.6
   SWE-Gym-7B (Pan et al., 2025) OpenHands SFT 10.6
   SWE-agent-LM-7B (Yang et al., 2025c) SWE-agent SFT 15.2
   Lingma-SWE-GPT-7B (Ma et al., 2025) SWESynInfer SFT 18.2
   SWE-Mirror-LM-7B (Wang et al., 2025b) MOpenHands SFT 22.8
   SWE-Dev-7B (Wang et al., 2025a) OpenHands RL 23.4
   Klear-Agent-8B-SFT (Kwai-Klear, 2025) Mini-SWE-agent-plus SFT 39.0
   SWE-Lego-Qwen3-8B OpenHands SFT 44.4 / 42.2
    + TTS@16 OpenHands SFT 53.4 / 49.6
Parameters == 32B
   Qwen2.5-Coder-Instruct-32B (Hui et al., 2024) MOpenHands - 6.2
   Qwen3-32B (Yang et al., 2025a) OpenHands - 23.2
   SWE-Gym-32B (Pan et al., 2025) OpenHands SFT 20.6
   R2E-Gym-32B (Jain et al., 2025) R2E-Gym SFT 34.4
    + TTS@16 R2E-Gym SFT 49.4
    + TTS@26 R2E-Gym SFT 51.0
   SWE-Dev-32B (Wang et al., 2025a) OpenHands RL 36.6
   Skywork-SWE-32B (Zeng et al., 2025) OpenHands SFT 38.0
    + TTS@8 OpenHands SFT 47.0
   SWE-agent-LM-32B (Yang et al., 2025c) SWE-agent SFT 40.2
   DeepSWE-32B-Preview (Luo et al., 2025) OpenHands RL 42.2
    + TTS@16 OpenHands RL 59.0
   SWE-Mirror-LM-32B (Wang et al., 2025b) MOpenHands SFT 52.2
   CWM-32B (Copet et al., 2025) Agentless CPT + SFT + RL 53.9
   SWE-Lego-Qwen3-32B OpenHands SFT 57.6 / 52.6
    + TTS@16 OpenHands SFT 60.4 / 58.8
Table 6: Performance comparison on the SWE-bench Verified. Our results are reported in the “A/B” format, representing the results with and without Git hacking respectively.
6 Related Work
6.1 Training Dataset for Coding
High-quality data is foundational for code intelligence. Early works focused on large-scale pretraining corpora like The Stack (Kocetkov et al., 2022), often employing rigorous filtering pipelines (Li et al., 2023) to ensure quality. The field has since shifted towards instruction tuning, utilizing self-instruct methods (Chaudhary, 2023), Git commit structures (Muennighoff et al., 2023), and synthetic generation techniques like Evol-Instruct (Luo et al., 2024) and OSS-Instruct (Wei et al., 2023). Broader instruction tuning efforts have also contributed significantly, with datasets like OpenHermes (Teknium, 2023) and Glaive (GlaiveAI, 2023) providing extensive general and code-specific training data. Recently, attention has turned to repository-level SWE datasets. Works such as R2E-Gym (Jain et al., 2025), SWE-smith (Yang et al., 2025c), and SWE-rebench (Badertdinov et al., 2025) provide scalable pipelines for task collection and environment construction. However, there remains a scarcity of large-scale, verifiable expert trajectories for long-horizon issue fixing, a gap our hybrid dataset addresses.
6.2 Coding LLMs and Agents
Recent progress spans both powerful foundation models and agentic frameworks that enable autonomous software engineering. Foundation models such as Code Llama (Rozière et al., 2023) and DeepSeek Coder (Guo et al., 2024) continue to serve as critical backbones, achieving state-of-the-art performance through massive code pretraining and extended context windows. To handle repository contexts, some approaches integrate RAG (Shrivastava et al., 2023) or AST-based retrieval (Zhang et al., 2024). Agentic frameworks such as Agentless (Xia et al., 2024), SWE-agent (Yang et al., 2024), Mini-SWE-Agent (Team, 2024b), Mini-SWE-Agent-Plus (Kwai-Klear, 2025) OpenHands (Team, 2024a) and MOpenHands (Zan et al., 2025) serve as scaffolds to streamline interactions with development environments. Others emphasize planning (Bairi et al., 2023) that views the repository-level coding as a planning problem or self-evolving the SWE agents on the fly(Xia et al., 2025).
6.3 Benchmarks for Software Engineering
Evaluation has progressed from function-level synthesis (Chen et al., 2021b) to repository-level tasks. Benchmarks like ClassEval (Du et al., 2023), CrossCodeEval (Ding et al., 2023), and RepoBench (Liu et al., 2024) assess broader context utilization. For autonomous agents, the focus is on realistic issue resolving, utilizing datasets like Defects4J (Just et al., 2014) and environments like InterCode (Yang et al., 2023). SWE-bench (Jimenez et al., 2024) has become the standard for GitHub issue resolution, inspiring variants for multilingual (Zan et al., 2025), multimodal (Yang et al., 2025b), and comprehensive agentic evaluation (Xu et al., 2025). Safety benchmarks like CyberSecEval (Bhatt et al., 2024) further ensure code security. Our work aligns with this trend, focusing on scaling capabilities for practical software development.
7 Conclusions
We present SWE-Lego, a framework consisting of hybrid executable data construction, refined SFT, and verifier-guided TTS, which achieves strong hack-free performance on SWE-bench Verified. The core driver of performance gains lies in scalable, verifiable hybrid data instances and expert trajectories. Meanwhile, error masking and curriculum learning stabilize the training process, and our proposed generative verifier, built on parallel test-time scaling, efficiently allocates inference computing resources to boost the performance further. Collectively, these strategies provide a reproducible paradigm for building robust agentic systems for software engineering issue resolving.
Appendix
Appendix A Additional Information on Data Construction
A.1 Real and Synthetic Instance Construction
Here we detail how we collect or generate the core fields, golden_patch, FAIL-to-PASS, PASS_TO_PASS, problem_statement and image_name of SWE-Lego data instances.
  • golden_patch: For real-world instances, the golden_patch corresponds to the merged PR diff, representing the human-authored fix. For synthetic instances, we inject a bug and then derive the golden patch by reverting this injection, i.e., the diff between the buggy branch and the original clean commit. Specifically, real tasks use a buggy pre-commit version of the repository in the sandbox, with the patch serving as the golden solution. For synthetic tasks, following SWE-smith[Yang et al., 2025c]), we place the non-buggy repository version inside the sandbox, while the buggy version resides in a remote Git branch; here, the patch field records the bug injection logic. In practice, before solving a synthetic task, one must check out the buggy branch in the sandbox. The golden solution is to revert the patch specified in the patch field.
  • FAIL-to-PASS and PASS_TO_PASS: In real data, we trace tests that are introduced or modified in the PR: tests that fail on the pre-merge commit but pass after applying the patch form the FAIL_TO_PASS set, while tests that pass in both states are labeled as PASS_TO_PASS. For synthetic data, we run the repository’s test suite before and after the injected bug. Newly failing tests define FAIL_TO_PASS, and unaffected tests form PASS_TO_PASS. This shared test-based view ensures that every instance, regardless of origin, is grounded in executable signals about both correctness and regression.
  • problem_statement: For real-world instances, the problem_statement is extracted from the GitHub issue linked to the PR, preserving the natural language description written by developers and users. For synthetic instances, we prompt an LLM with the golden patch, failing tests, and execution logs to synthesize a realistic issue description that mimics how a human would report the bug. The result is a unified textual interface where real issues contribute authentic language, and synthetic issues expand coverage in a controlled manner.
  • image_name: Each real-world instance is associated with a unique Docker image (a.k.a., sandbox) built from the exact pre-merge commit, since different PRs can depend on distinct dependency states or build configurations. In contrast, synthetic instances from the same repository typically share a sandbox image based on a stable base commit, because bug injections operate at the code level without introducing new system-level dependencies.
A.2 Analysis of Data Categories
Category Description Real Data (%) Synthetic Data (%)
Top-5 Categories
A API / signature mismatch 26.6 12.8
B Logic / conditional bug 29.9 49.4
C Input validation / boundary / sentinel handling error 16.1 5.9
D Constructor / inheritance / inheritance contract break 2.5 7.5
E Missing import / symbol / attribute error 8.3 19.3
The Remaining Categories
F State consistency / bookkeeping / caching bug 6.4 2.6
G Copy semantics / mutability / in-place mutation of inputs 0.7 0.3
H Protocol / spec conformance bug 7.2 1.4
I IO / file system / resource handling bug 1.9 0.7
J Security / sensitive data leakage 0.4 0.1
Table 7: Distribution of categories in real-world and synthetic training dataset.
Following the categorization method of SWE instances proposed in BugPilot [Sonwane et al., 2025], we prompt the Claude-4-Sonnet to categorize the training instances SWE-Lego for analyses.
The comparison on Table 7 reveals a significant complementarity between the real and synthetic data sources in SWE-Lego. Real-world data tends to focus on “external interaction and boundary defense” (e.g., API changes, input validation, protocol conformance), whereas synthetic data emphasizes “internal implementation and structural integrity” (e.g., logic flow, dependency management). Beyond the inherent scalability of synthetic data generation, incorporating it into the training set offers distributional benefits by supplementing rare structural defects and rebalancing the focus between interface adaptation and internal logic.
Appendix B Additional Analysis on Test-Time Scaling
We present additional experiments on test-time scaling, analyzing the saturation in sequential scaling and the effects of training-data scale and verifier-model scale on parallel scaling.
B.1 Sequential Scaling Analysis
Truncation Rate and Average Turns.
As the turn limit increases for SWE-Lego-Qwen3-8B rollouts, the truncation rate declines quickly but becomes marginal beyond roughly 140 turns (Figure 11(a)). The average number of turns per trajectory similarly rises and then plateaus. Most trajectories either terminate (success or failure) or fall into unproductive states well before exhausting large turn budgets, so additional turns beyond this threshold yield little marginal utility.
Resolve Rate by Turn Count.
Resolve rates are low for very small turn counts (insufficient exploration) and also for very large counts (the agent is typically stuck, retrying ineffective approaches), with the highest rates appearing at intermediate ranges (Figure 11(b)).
These observations provide empirical support for the sequential scaling saturation phenomenon: beyond a certain turn limit, additional sequential compute yields limited marginal benefit. This motivates the shift to parallel scaling with verifier selection once sequential scaling saturates.
(a) Truncation Rate and Average Turns
(b) Resolve Rate by Turn Count
Figure 11: Analysis of sequential scaling behavior. (a) Truncation rate and average turns under different maximum turn limits: both metrics show diminishing changes beyond 140 turns. (b) Resolve rate across turn count intervals: trajectories with very few or very many turns tend to have lower resolve rates.
B.2 Impact of Training Data Scale and Model Scale
Training Data Scale.
We ablate verifier training data volume on SWE-Lego-Qwen3-8B rollouts by comparing the full set (18K trajectories: 5K resolved + 13K unresolved) to a 6K subset matching the scale used by R2E-Gym-Verifier. As shown in Figure 12(a), the 18K-trained verifier consistently outperforms the 6K verifier across all $K$. At TTS@16, it achieves 49.6% vs. 47.6%, with the gap widening as $K$ increases, indicating that larger training sets improve discrimination in larger candidate pools. Notably, with the same 6K scale, our verifier slightly outperforms R2E-Gym-Verifier-14B (47.6% vs. 47.0%), suggesting that data quality and alignment with the rollout agent also matter, beyond sheer quantity.
Verifier Model Scale.
We compare 8B and 30B verifiers (Qwen3-8B and Qwen3-Coder-30B-A3B backbones) when scoring rollouts from SWE-Lego-Qwen3-8B and SWE-Lego-Qwen3-32B agents (Figure 12(b)). For 8B rollouts, 8B and 30B verifiers perform similarly across $K$. For 32B rollouts, the 30B verifier is consistently better, especially at larger $K$; at TTS@16, it reaches 58.8% vs. 56.6%. One possible e explanation is that 32B outputs contain subtler errors that benefit from larger verifier capacity; as $K$ grows, more near-correct candidates increase the premium on fine-grained ranking.
(a) Impact of Training Data Scale
(b) Impact of Verifier Model Scale
Figure 12: Ablation studies on verifier training. (a) Training data scale: verifiers trained on 18K trajectories outperform those trained on 6K trajectories. (b) Model scale: for 8B rollouts, verifier size has minimal impact; for 32B rollouts, the 30B verifier provides gains over the 8B verifier, especially at higher $K$.
Appendix C Extensions and Future Work
Beyond Issue Resolving.
Extending the scope beyond defect repair to other tasks, like feature implementation, refactoring or tests generation, reveal limitations in current training data and benchmarks. Constructing datasets that adequately encode ambiguous requirements, nuanced acceptance criteria, and non-regression constraints remains challenging. Moreover, validating agent performance across richer dimensions, such as functional correctness, performance bounds, and interface stability, requires more complex evaluation frameworks than those used for bug fixing. Supervising higher-order behaviors like design decisions, system integration, and code refactoring also lacks well-defined signals for effective learning.
Multilingual SWE.
Expanding beyond Python to languages such as JavaScript/TypeScript, Java, Go, and Rust is impeded by language-specific toolchains, build systems, and testing frameworks. The absence of cross-language sandboxes and multilingual datasets limits agents’ ability to generalize across ecosystems. Further challenges include navigating heterogeneous dependency management practices, adapting to diverse testing paradigms, and accounting for language-specific code organization idioms, all of which complicate the transfer of core agentic skills (navigation, localization, editing, validation) across programming languages.
Cross-task Capability Balance.
Heterogeneous code tasks, including repository-level issue resolution, file-specific code reasoning, and competitive coding, exhibit divergent demands on agent capabilities, leading to risks of uneven performance. Balancing proficiency across these tasks is hindered by varying difficulty gradients and interaction requirements, which may cause agents to overfit to dominant task archetypes. The lack of standardized metrics for comparing capability across such distinct tasks further complicates efforts to assess and mitigate this imbalance.
Appendix D Example of Git Hacking
In Figure 13, we present a concrete example of Git hacking: an SWE agent peeks at the Git commit history and unintentionally retrieves the pre-existing solution. During evaluation, the agent may run common Git commands (e.g., git log, git diff) to bypass independent problem-solving and directly replicate the golden patch, falsely inflating performance and invalidating the evaluation of genuine software engineering capabilities.
Figure 13: Demonstration of Git hacking by SWE agents.